Assuming the system is well insulated, what temperature would 490g of water reach, starting at 15∘C, if a 60W heater heated it for 20minutes? [Specific heat capacity of water=4200J/(kgK)]
1
Expert's answer
2018-06-06T10:42:08-0400
The heat will be
Q=mc(T-T_0 )
T=T_0+Q/mc=T_0+Pt/mc
T=15+(60)(20∙60)/(0.49)(4200) =50℃.